OpenStudy (gabby1214):

Need help!!! 1.When writing linear equations how do you determine which form of line to use? 2.Write an equation in point slope form the line that passes through the following point, then use the same set of points to write the equation in standard form and again in slope intercept form set of points: (0,-2), (3,2)

jhonyy9 (jhonyy9):

do you know what mean ,,linear equation" ?

OpenStudy (trojanpoem):

standard form : Ax + By = C (A, B and C are constants) Slope-intercept form: y = mx + b (m: slope, b : y-intercept)

OpenStudy (anonymous):

1. You would use different equations for what you want to use the equation for. You would use slope intercept form for graphing, for example. 2. This is easy with y-y= m(x-x) aka point slope form. Y-2= 4/3(X-0) Y-2= 4/3 X Y= 4/3X +2. Tell me if you need any more clarification.

OpenStudy (anonymous):

Here's it in standard form: 4/3x+y = 2

OpenStudy (danjs):

by the information you know about the line or that was given to you the second one demonstrates that i think.. the amount of work to put together each of the forms, given 2 points